OpenStudy (anonymous):

a machine in a soft drink factory fills 840 in 6 hours. how many can it fill in 5 hours

OpenStudy (anonymous):

700

OpenStudy (saifoo.khan):

hours : Liters 6 : 840 5 : x Cross multiply

OpenStudy (anonymous):

but that leaves u with 840 still @seafood

OpenStudy (saifoo.khan):

No it dosnt! :P

OpenStudy (anonymous):

oh..lol I STILL GOT IT RITE1!!!

OpenStudy (saifoo.khan):

mine is right as well

OpenStudy (saifoo.khan):

\[6x = 840 \times 5\]

OpenStudy (anonymous):

KK...i gues i saw it wrong..srry smthing got in ma eye

OpenStudy (saifoo.khan):

LOL. a bug might be

OpenStudy (anonymous):

u sure like offend pple huh

OpenStudy (anonymous):

um whats the answer

OpenStudy (anonymous):

oh srry, its 700

OpenStudy (anonymous):

how did yu get it

OpenStudy (saifoo.khan):

i told u how to

OpenStudy (anonymous):

cross multiply

OpenStudy (anonymous):

ORRR

OpenStudy (anonymous):

6x=5(840)

OpenStudy (anonymous):

uh huh i tried it myself i got 1048 yall are the genuis i would really need a better understanding

OpenStudy (anonymous):

6x=4200

OpenStudy (anonymous):

divide by 6

OpenStudy (anonymous):

leaving you with 700

OpenStudy (anonymous):

do you understand it?

OpenStudy (anonymous):

oh srry my mistake i multiply 840 by 6

OpenStudy (anonymous):

lol i did that the first time too...

OpenStudy (anonymous):

yes thanks alot

OpenStudy (anonymous):

no problemo

OpenStudy (saifoo.khan):

problem*

OpenStudy (anonymous):

oh my gosh....

OpenStudy (anonymous):

ill definetely get an a at test tommorrow what saifoo???

OpenStudy (saifoo.khan):

Nothing. :p

OpenStudy (saifoo.khan):

i hope u get an A+.

OpenStudy (anonymous):

yep im sure u would, i gotta feeling im a get an F on ma alg 2 test tmmrw....

OpenStudy (saifoo.khan):

Alg 2 is easy! :P

OpenStudy (saifoo.khan):

F for fantanstic?

OpenStudy (anonymous):

ahh shut up!! ur makin it worse seafood

OpenStudy (saifoo.khan):

LOOOOOOOOOOOOOL

OpenStudy (anonymous):

lol crazy people what is alg? algebra?

OpenStudy (anonymous):

yep, and saifodeen isnt helpin here

OpenStudy (saifoo.khan):

haha, Dont worry, if you get a F.. so think f for fantastic.

OpenStudy (anonymous):

oh it not hard just focus on finding your missing number think of it as finding your missing bf of gf :p

OpenStudy (saifoo.khan):

Haha.

OpenStudy (anonymous):

lol tats ur stragety? i gotta make formulas into songs for me to member them...

OpenStudy (saifoo.khan):

ex-bf as the questions says find "x"

OpenStudy (anonymous):

wats ur stragety seafood

OpenStudy (saifoo.khan):

to make it as simpler as possible and go for it.

OpenStudy (anonymous):

i dunt get it

OpenStudy (anonymous):

i got a bad case of test anxiety

OpenStudy (saifoo.khan):

lol.

OpenStudy (anonymous):

no mines is looking at others paper lol jking ill explan an example 6x2+m=`19 12-12+m=19-12(minus at both sides) M=7 i found seven missing boyfrends

OpenStudy (anonymous):

thats not funnay, i always fail a test and the second i get it back i know all ma small stupid mistakes....

OpenStudy (saifoo.khan):

m = boyfriends? o.O

OpenStudy (anonymous):

lol i love lovelybabes stragety...its so funnay

OpenStudy (anonymous):

m= mom best frend brother(boyfrend )lol

OpenStudy (anonymous):

wat????

OpenStudy (saifoo.khan):

DAMN LOL.

OpenStudy (anonymous):

m=ur moms best friends brother??? this is confusing me.....

OpenStudy (saifoo.khan):

lol, you are X-man. u should understand that as fast as possible.

OpenStudy (anonymous):

lollllllllllllll joking you all crazy up in here

OpenStudy (anonymous):

heyy...saifodeen..tat not rite, since wen does x-man have math abillities

OpenStudy (anonymous):

goodnight before my parent catch me lol

OpenStudy (anonymous):

ur going to sleep? this early?

OpenStudy (saifoo.khan):

haha, then which "man" have that abilities?

OpenStudy (saifoo.khan):

those*

OpenStudy (anonymous):

uhhh..i dunno u think of them

OpenStudy (saifoo.khan):

Sorry dont have enough time.

OpenStudy (anonymous):

ya me too...gotta go

OpenStudy (saifoo.khan):

everybdy is going! :/

OpenStudy (anonymous):

then why arent u guys leaving

OpenStudy (saifoo.khan):

nobdy should leave.

OpenStudy (anonymous):

im still here

OpenStudy (saifoo.khan):

:DDDD

OpenStudy (anonymous):

oh...ur her too..wat luck

OpenStudy (saifoo.khan):

her? :O what her? who her?

OpenStudy (anonymous):

im meant here!!!

OpenStudy (saifoo.khan):

Oh yes Luckk... badman

OpenStudy (anonymous):

wat ever saifodeen

OpenStudy (anonymous):

ha how do u like ur new names

OpenStudy (saifoo.khan):

Yea, whatever, where ever Spiderman.

OpenStudy (saifoo.khan):

My names are fine... xDD

OpenStudy (anonymous):

sure seafood

OpenStudy (saifoo.khan):

;)